Calculate the elasticity of supply if a 15 %increase in the price of soya bean oil increases 

its supply from 300 to 345 units


1
Expert's answer
2021-05-17T17:29:02-0400

es = % change in supply / % change in price

change in supply = (100 x (Q1-Q)) / Q = (100 x (345-300)) / 300 = 15%

es = 15 / 15 = 1